# Deep Learning based discovery of Integrable Systems

### v1

@@ Abstract
We introduce a novel machine learning based framework for discovering integrable models. Our approach first employs a synchronized ensemble of neural networks to find high-precision numerical solution to the Yang-Baxter equation within a specified class. Then, using an auxiliary system of algebraic equations, $[Q_{2},Q_{3}]=0$, and the numerical value of the Hamiltonian obtained via deep learning as a seed, we reconstruct the entire Hamiltonian family, forming an algebraic variety. We illustrate our presentation with three- and four-dimensional spin chains of difference form with local interactions. Remarkably, all discovered Hamiltonian families form rational varieties.
@@ I Overview
    Hamiltonian Extraction: We look for an algebraic variety of integrable models in the vicinity of the numerically found Hamiltonian. The general way to do this is somewhat involved and we postpone the discussion to Section V. One point worth mentioning here is that the method for extracting exact relations relies on the integer nature of the coefficients in the polynomial relations between Hamiltonian entries, which enables rounding of numerical data. While we do not have a formal proof yet, our findings thus far suggest that deriving polynomial relations with integer coefficients has been sufficient for solving 16. One of the simplest integrable families we discovered from the (1) pattern is a 25-vertex model that forms the following three-parameter linear variety.:
@@
    $R\left(u\right)=P\exp\left(u\,H\right)\,,$ (3)
@@ IV.1 Averting Mode Collapse
where $\mu_{k},\lambda_{repul}$ are further new hyperparameters. In practice, we fix these hyperparameters randomly to some positive values between 0 and 1 at the start of each optimization run. For example if we want to repulse from the given family of Hamiltonians we can choose $\{C_{k}\}$ as a Gröbner basis $G_{k}(\{h_{ij}\})$ of the ideal defining this subvariety and set $\lambda_{repul}=0$, $\mu_{k}=1$.
@@ V Hamiltonian Extraction
In all cases we analyzed, the coefficients $\tilde{c}_{\beta,i}$ were very close to integer numbers if the precision for $\mathcal{V}$ was high enough, which suggests that we should just round them $c_{\beta,i}=Round[\tilde{c}_{\beta,i}]$ and use
@@
The crucial element of this procedure is the integrality of coefficients $c_{\beta,i}$. Although a formal proof is not yet available, all cases studied so far indicate that extracting polynomial relations with integer coefficients has been sufficient to obtain a solution.
@@ VII Discussion and future directions
From the algebraic geometry point of view we construct irreducible components of the algebraic variety defined by (16). The extraction of exact relations described in section V crucially relies on the fact that coefficients are integers, which allows us to round off our numerical data. While we do not yet have a formal proof, all our cases so far have demonstrated that extracting polynomial relations with integer coefficients has been sufficient to solve 16. Moreover all integrable Hamiltonians we found, remarkably, can be written as rational varieties and we would like to conjecture that all irreducible components (18) are (uni)rational varieties. It is possible in principle, that the conjecture may eventually need to be qualified by an as yet unknown, but fairly general extra constraint. Reshetikhin condition has the form of the vanishing commutator, and in the simpler case of commuting varieties the rationality was already proved [34]. Some of the discovered Hamiltonians form highly complex varieties – see e.g. (46) – making it tempting to initiate their systematic analysis using algebraic geometry. In particular, studying their topology by computing homology groups could reveal distinctive properties of this class of varieties and inspire a new research direction in the algebro-geometric analysis of integrable models.

### v2

@@ Abstract
We introduce a novel machine learning based framework for discovering integrable models. Our approach first employs a synchronized ensemble of neural networks to find high-precision numerical solution to the Yang-Baxter equation within a specified class. Then, using an auxiliary system of algebraic equations, $[Q_{2},Q_{3}]=0$, and the numerical value of the Hamiltonian obtained via deep learning as a seed, we reconstruct the entire Hamiltonian family, forming an algebraic variety. We illustrate our presentation with three- and four-dimensional spin chains of difference form. Remarkably, all discovered Hamiltonian families form rational varieties.
@@ I Overview
    Hamiltonian Extraction: We look for an algebraic variety of integrable models in the vicinity of the numerically found Hamiltonian. The general way to do this is somewhat involved and we postpone the discussion to Section V. One point worth mentioning here is that the method for extracting exact relations relies on the integer nature of the coefficients in the polynomial relations between Hamiltonian entries, which enables rounding of numerical data. While we do not have a formal proof yet, our findings thus far suggest that deriving polynomial relations with integer coefficients has been sufficient for solving 16. One of the simplest integrable families we discovered from the (1) pattern is a 25-vertex model that forms the following three-parameter linear variety $H_{25}$:
@@
    $R\left(u\right)=P\exp\left(u\,H_{25}\right)\,,$ (3)
@@ IV.1 Averting Mode Collapse
where $\mu_{k},\lambda_{repul}$ are further new hyperparameters. In practice, we fix these hyperparameters randomly to some positive values between 0 and 1 at the start of each optimization run. For example if we want to repulse from the given family of Hamiltonians we can choose $\{C_{k}\}$ as a Gröbner basis $G_{k}(\{h_{ij}\})$ of the ideal defining this subvariety and set $\lambda_{repul}=1$, $\mu_{k}=1$.
@@ V Hamiltonian Extraction
In all cases we analyzed, the coefficients $\tilde{c}_{\beta,i}$ were very close to either integers or rational numbers with small numerator and denominator if the precision for $\mathcal{V}$ was high enough, which suggests that we should just round them $c_{\beta,i}=Round_{\mathbb{Q}}[\tilde{c}_{\beta,i}]$ to the ”closest” rational number and use
@@
The crucial element of this procedure is the rationality of coefficients $c_{\beta,i}$. Although a formal proof is not yet available, all cases studied so far indicate that extracting polynomial relations with integer coefficients111after the multiplication by the common denominator we can speak just about integer numbers instead of rational has been sufficient to obtain a solution.
@@ VII Discussion and future directions
From the algebraic geometry point of view we construct irreducible components of the algebraic variety defined by (16). The extraction of exact relations described in section V crucially relies on the fact that coefficients are integers, which allows us to round off our numerical data. While we do not yet have a formal proof, all our cases so far have demonstrated that extracting polynomial relations with integer coefficients has been sufficient to solve 16. Moreover all integrable Hamiltonians we found, remarkably, can be written as rational varieties and we would like to conjecture that all irreducible components (18) are (uni)rational varieties. It is possible in principle, that the conjecture may eventually need to be qualified by an as yet unknown, but fairly general extra constraint. Reshetikhin condition has the form of the vanishing commutator, and in the simpler case of commuting varieties the rationality was already proved [35]  [^2] We thank Samuel Stark for bringing this reference to our attention.. Some of the discovered Hamiltonians form highly complex varieties – see e.g. (46) – making it tempting to initiate their systematic analysis using algebraic geometry. In particular, studying their topology by computing homology groups could reveal distinctive properties of this class of varieties and inspire a new research direction in the algebro-geometric analysis of integrable models.